OpenStudy (anonymous):

SOMEONE PLEASE HELP ME!! I'll give a medal! The graph below shows a system of equations: The x-coordinate of the solution to the system of equations is __________.

OpenStudy (whpalmer4):

You know that the solution to the system of equations is the point where all the lines intersect, right?

OpenStudy (anonymous):

Uhm no haha. Math is my worst subject, I am surprised I even moved on to Algebra.

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

Wait so it's 1?

OpenStudy (anonymous):

Never mind! I was looking at the wrong graph!!

OpenStudy (whpalmer4):

Okay, well, each of the straight lines in the graph represents the set of solutions to one of the equation. The solution to a system of equations is going to be where those lines cross, because they cross at a point they share, right?

OpenStudy (whpalmer4):

Is the graph you posted the link to the correct graph?

OpenStudy (anonymous):

Well yes, they meet at 2 right? And yes that's the correct graph

OpenStudy (anonymous):

Either 2 or -2?

OpenStudy (whpalmer4):

the x coordinate (the value on the x-axis) is the the right-left value, and the y coordinate (the value on the y-axis) is the up-down value. given that, which do you think is correct?

OpenStudy (whpalmer4):

and I can guarantee that the person who made this problem chose to have one be 2 and the other be -2 just to confuse you (well, to see how well you understand, really)!

OpenStudy (anonymous):

So 2? Because it says x-coordinate

OpenStudy (anonymous):

-2 I meant

OpenStudy (whpalmer4):

okay, that's better :-)

OpenStudy (anonymous):

Okay thank you! Could you help me with a few other please??

OpenStudy (anonymous):

others*

OpenStudy (whpalmer4):

lol — do you find many guys that say no to you? :-)

OpenStudy (anonymous):

Well a lot have lol, I hope this means you'll help?:)

OpenStudy (whpalmer4):

of course, that's why I'm here instead of fixing the dishwasher :-)

OpenStudy (whpalmer4):

you'll fix my dishwasher in exchange, right? :-)

OpenStudy (anonymous):

If you need to do that I can wait haha it's okay, as long as you come back! lol

OpenStudy (anonymous):

I can't guarantee I'll fix it but all I can do is try right? :)

OpenStudy (whpalmer4):

okay, let's see the next problem. my goal is to be so successful in showing you how to do them that you won't need to ask many!

OpenStudy (anonymous):

Sally bought x pounds of apples and y pounds of pears. Three times the number of pounds of pears she bought was 3 more than the number of pounds of apples she bought. The total number of pounds of apples and pears she bought was 5 pounds. Which graph best shows the number of pounds of apples and pears Sally bought? https://gbwebacademy.brainhoney.com/Resource/22373151,A49,0,38/Assets/73504_53bff066/module5_graph85.gif https://gbwebacademy.brainhoney.com/Resource/22373151,A49,0,38/Assets/73504_53bff066/module5_graph89.gif https://gbwebacademy.brainhoney.com/Resource/22373151,A49,0,38/Assets/73504_53bff066/module5_graph86.gif https://gbwebacademy.brainhoney.com/Resource/22373151,A49,0,38/Assets/73504_53bff066/module5_graph87.gif

OpenStudy (whpalmer4):

Okay, this one is a bit trickier. You'll need to write some equations to represent what the words in the problem say. Sally bought x pounds of apples and y pounds of pears. [...] The total number of pounds of apples and pears she bought was 5 pounds. Can you write an equation that says that?

OpenStudy (anonymous):

x+y=5?

OpenStudy (whpalmer4):

excellent!

OpenStudy (whpalmer4):

"Three times the number of pounds of pears she bought was 3 more than the number of pounds of apples she bought." Now, how about this one?

OpenStudy (anonymous):

Well this one is confusing me a little bit... Would this make sense for pears? 3x=y Probably not but I am trying..

OpenStudy (whpalmer4):

well, just go through the sentence replacing the words with the variable names apples are x pears are y so three times y is 3 more than x what is that in equation form?

OpenStudy (whpalmer4):

"three times y" becomes \(3y\) "is" becomes \(=\) "3 more than x" becomes \(3+x\) so that gives us...

OpenStudy (anonymous):

3y=3+x Sorry I was getting there, I know I am slow/:

OpenStudy (anonymous):

is it suppose to be 3y=x+3 or is it fine the way it is?

OpenStudy (whpalmer4):

Okay, so we have two equations, and that means we have two lines (in theory, they could end up being the same line, but I can tell by looking that they are not, and so can you!) As we said before, the solution of the system of equations will be the point where all of the lines intersect. \[3y=x+3\]is just as good as \[x+3 = 3y\]or\[x = 3y-3\]or any of the other arrangements

OpenStudy (whpalmer4):

Now we need to figure out which of those graphs correspond to our equations. One way would be to plug a few numbers into the equation, and see if the results appear on the graph we are considering. I always like to try 0 because it usually makes the work easy. So let's plug 0 into both of our equations for \(x\) and see what pops out for \(y\). When we do so, we'll have a pair of points, and we can look at each graph and discard any graph which does not have both of those points on the lines.

OpenStudy (whpalmer4):

\[x+y=5\]well \(x=0\) so \[0+y=5\]means that \[y=5\]so our first point is \((0,5)\) any question about what I did there?

OpenStudy (anonymous):

Uhm, I don't think so. I understand.

OpenStudy (anonymous):

So far...

OpenStudy (whpalmer4):

okay, can you plug \(x=0\) in to the other equation and find out the corresponding value for \(y\)?

OpenStudy (anonymous):

Well where are you getting the equations? From the graphs?

OpenStudy (whpalmer4):

no, we just figured them out from the story problem, remember? the other equation is \[3y = 3+x\]

OpenStudy (whpalmer4):

now, if we were serious math geeks, we could look at the graphs and say "oh, the equation for this line is such-and-such" and work from that direction, but we aren't, so we won't :-)

OpenStudy (anonymous):

I don't want to look stupid and give the wrong answer... /:

OpenStudy (whpalmer4):

nah, you won't look stupid. if it looks like you gave the wrong answer, I'll avert my eyes quickly :-)

OpenStudy (whpalmer4):

so we substitute \(0\) for \(x\) in the equation \[3y = 3+x\]which gives us \[3y = 3 + 0\] or \[3y = 3\]

OpenStudy (anonymous):

y+3=x x=0 so y+3=0 so y=3? Don't laugh.. I'm sure i look really dumb. I was just looking at what you did for the first one..

OpenStudy (whpalmer4):

mmm...not quite. our first equation was \[x+y = 5\]and after I replaced \(x\) with \(0\) that became \[0+y=5\]\[y=5\]so our first point is \((0,5)\) Our second equation was \[3y=3+x\] and after I replaced \(x\) with \(0\) that became \[3y = 3 + 0 \]\[3y=3\]divide both sides by the coefficient of \(y\) and we get\[\frac{3y}{3} = \frac{3}{3}\]\[y = 1\]so our second point is \((0,1)\) clear as mud?

OpenStudy (anonymous):

Oh okay, well that seemed really easy. Now I feel stupid...

OpenStudy (whpalmer4):

Now, which of the graphs have both \((0,5)\) and \((0,1)\) as points on the lines?

OpenStudy (whpalmer4):

Yes, taking \(x=0\) makes the arithmetic easy most of the time. Sometimes \(x=1\) is also easy, if you need to repeat the process at a different spot. \(x=539.427634\) is usually NOT easy :-)

OpenStudy (whpalmer4):

the "easy" values are also easy to find on the graph

OpenStudy (anonymous):

Graph b?

OpenStudy (whpalmer4):

If graph b is this one, then yes:

OpenStudy (anonymous):

Yes it is :)

OpenStudy (whpalmer4):

As a check, notice that the lines intersect at a point - can you tell me the x-coordinate of that point?

OpenStudy (anonymous):

3?

OpenStudy (whpalmer4):

Very good. To double-check our choice, let's figure out from the equations what the values of \(y\) should be at \(x=3\). If we did everything correctly, the values will be the same as the \(y\) coordinate of that point where they intersect.

OpenStudy (anonymous):

Wow, I didn't realize how easily you can figure this stuff out. Thank you so much! Mind helping with a few more? If you have the time...

OpenStudy (whpalmer4):

There's a saying in medical school: "see one, do one, teach one" You saw how I found the \(y\) coordinates for \(x=0\) Now you get to find the \(y\) coordinates for \(x=3\) Then you'll be all set to teach a friend :-)

OpenStudy (whpalmer4):

\[x+y=5\]\[3+y=5\]\[3-3+y=5-3\]\[y=2\]and\[3y=3+x\]\[3y=3+3\]\[3y=6\]\[y=2\]so both of our equations supply \((3,2)\) as a point on the graph, and our chosen graph does have \((3,2)\) as a point of intersection, so we did it correctly. Next problem?

OpenStudy (anonymous):

The graph of the following system of equations is 6x + 3y = 12 A. overlapping lines B. parallel lines C. intersecting lines

OpenStudy (anonymous):

and 2x + y = 6

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

okay, our equations are \[6x+3y=12\]\[2x+y=6\] do you notice anything about the coefficients of the two equations?

OpenStudy (anonymous):

Uhm no not necessarily..

OpenStudy (whpalmer4):

sorry, had an interruption compare the coefficients of the \(x\)'s now compare the coefficients of the \(y\)'s see a pattern?

OpenStudy (anonymous):

oh no worries, and i think so.

OpenStudy (whpalmer4):

the coefficient in the first equation is 3x the coefficient in the second equation, right?

OpenStudy (anonymous):

Wait would the second one be 6?

OpenStudy (whpalmer4):

if we have an equation of a line, say \[ax + by = c\]where \(a,b,c\) are constant values and \(x,y\) are variables, we can multiply each term of the equation by the same value and the resulting equation draws exactly the same line. For example, \[x+y=5\] from our previous problem. If we multiply the equation by 2, we get \[2x + 2y = 10\] Let's find out what the points would be on that line at \(x=0\) and \(x=3\): \[2(0)+2y=10\]\[0+2y=10\]\[2y=10\]\[y=5\]so \((0,5)\) is a point again \[2(3)+2y=10\]\[6+2y=10\]\[2y=4\]\[y=2\]so \((3,2)\) is also a point again If the equation of one line is a multiple of the equation of another, the two lines are identical, or coincident.

OpenStudy (anonymous):

Hmm, this is getting confusing.

OpenStudy (whpalmer4):

However, if they have different values for the constant term, but the coefficients are multiples like this, then they are parallel. \[x+y=5\]\[x+y=1\]are parallel lines, because for any given value of \(x\), the \(y\) value for one is going to be exactly \(5-1=4\) larger than the \(y\) value for the other

OpenStudy (whpalmer4):

some pictures might help! can you wait a few minutes, I have to do something for She Who Must Be Obeyed

OpenStudy (anonymous):

Haha yes that's completely fine with me

OpenStudy (whpalmer4):

Here is a plot of \[x+y=5\]

OpenStudy (whpalmer4):

and here is a plot of \[x+y=1\]

OpenStudy (whpalmer4):

Here is a plot of both of them on the same piece of paper:

OpenStudy (whpalmer4):

Those are parallel lines, agreed?

OpenStudy (anonymous):

Agreed

OpenStudy (whpalmer4):

Because they are parallel, they will not intersect. For any value of \(x\) we plug into \[x+y=1\]the value of \(y\) we get will be 4 less than the value of \(y\) we get plugging the same value of \(x\) into \[x+y=5\]

OpenStudy (whpalmer4):

Now I was talking about multiplying an equation by a constant not affecting the line at all. Let's do a picture of that. Here is \[x+y=5\] again

OpenStudy (anonymous):

Okay I am slowly understanding

OpenStudy (whpalmer4):

Here is \[2x+2y=10\]

OpenStudy (whpalmer4):

Exactly the same!

OpenStudy (anonymous):

Oh yeah it is the same

OpenStudy (whpalmer4):

So, if we can make one equation into an exactly copy of the other equation by just multiplying (or dividing) by a number, the two lines are coincident (or overlapping as your answer key puts it). We can't do that with this problem because while multiplying \[3*2x+3*y=3*6\] gives us \[6x+3y\]on the left hand side of the equals sign, which is exactly what our other equation has, the value on the right hand is still different. That means that the lines are parallel, but do not overlap. The third case is if the lines intersect in a point. For this to happen, the coefficients cannot be a multiple of each other. If they are not, somewhere on the infinite plane that contains the two lines, they will intersect. If you arrange the equation of a line in the form \[Ax+By=C\] where \(A,B,C\) are all constants (numbers), you can pretty quickly judge which of the three cases you have. If multiplying one equation by any number will give you matching values of \(A,B,C\), the two lines are overlapping. This means there are infinitely many solutions that satisfy the system of equations. If multiplying one equation by any number will give you matching values of \(A,B\) but not \(C\), then the lines are parallel. This means there are 0 solutions that satisfy the system of equations. Finally, if you can't multiply by any number to get a match of \(A,B\) for the two equations, that means the lines intersect. There will be 1 solution that satisfies the system of equations, and that is the poin

OpenStudy (whpalmer4):

is the point of intersection.

OpenStudy (whpalmer4):

I'm sure your eyes must have glazed over by now, right? :-)

OpenStudy (anonymous):

That's just a lot to read and understand but I think if I read it over closely I will understand but from where I am looking, they don't overlap but they are parallel? but you said something is the point of intersection but I am not sure what you were talking about

OpenStudy (whpalmer4):

Okay, quick recap: our equations were \[6x+3y=12\]\[ 2x+y=6\] If we multiply the second equation by 3, the coefficients for the \(x\) and \(y\) terms match each other: \[6x+3y=12\]\[ 3*2x+3*y=3*6\] \[6x+3y=12\]\[ 6x+3y=18\] but the constant on the other side of the equals sign is different. That means we have parallel lines. By my classification scheme, we have matching \(A,B\) but not \(C\)

OpenStudy (whpalmer4):

If we plot them both, here's what we get:

OpenStudy (anonymous):

So they are parallel

OpenStudy (whpalmer4):

Yep, that's right!

OpenStudy (whpalmer4):

We covered more ground than necessary to solve the problem, but the concepts are important, and my hope was that getting a good handle on them will help you with many problems.